Slope – 2/3, y – intercept (0, 5)
Answer by Cromlix(4381)

Using equation:
y = mx + c
y = 2/3x + 5
I'm not sure as to whether the slope
is -2/3 or 2/3
I have used 2/3.
If the slope is -2/3
the equation would be:
y = -2/3x + 5.
Hope this helps.
:-)